OpenStudy (anonymous):

solve and check -6x - 24 = 0

OpenStudy (anonymous):

-6x - 24 = 0 i.e. -6x = 24 i.e. x = 24/-6 i.e. x = -4 @Eagle8F000

OpenStudy (anonymous):

-6x=24 x=24/6 x=-4

OpenStudy (jamanorthi):

its gonna be - 4

OpenStudy (anonymous):

Thank You all! I see my mistake! Thanks so much! Thanks for tracing it out for me madrockz! You Rock!

OpenStudy (anonymous):

most welcome!!!!

OpenStudy (anonymous):

Ooops!! Sorry dpasingh! You Rock too! :-)